# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v2

@@ 5.2 Prompting to Ignore Misleading Comments
Table 7: Impact of explicit comment-ignoring prompts under MDC and MHC perturbations.
Model Inference VAN MDC MHC
@@
DeepSeek-V3 Direct $67.78$ $39.74$ $41.75$ $47.67$ $47.81$
CoT $92.07$ $70.15$ $82.05$ $89.98$ $94.57$
GPT-4o Direct $64.51$ $45.72$ $45.93$ $47.32$ $46.76$
CoT $94.36$ $87.47$ $89.35$ $94.78$ $94.57$
Qwen2.5-72B-Instruct Direct $54.91$ $47.95$ $48.02$ $32.64$ $32.50$
CoT $90.40$ $85.39$ $86.01$ $82.67$ $84.34$
@@
Next, we conduct a controlled experiment by instructing models: “Please ignore the comments and any other non-code elements in the code snippet.” Table 7 summarizes Pass@1 results on LCB for three high-performing models under both direct and CoT inference modes. Despite explicit instructions to ignore comments, the Pass@1 accuracy remains largely unaffected under direct inference, supporting our findings from §4.2.1 about LLMs’ inherent reliance on embedded NL cues. Although we observe improvements after employing CoT prompting, the Pass@1 performance under MDC is still substantially below vanilla settings. These results indicate that even with step-by-step execution, LLMs cannot fully detach from misleading textual distractions, highlighting a fundamental limitation in their reasoning controllability and dependence on NL cues.
5.3 Revealing the Vulnerability of Self-reflection Mechanisms
Figure 2: Quadratic increase ($R^{2}=0.9991$) in confusion tokens (e.g., “hmm”, “wait”, “perhaps”) during QwQ-32B’s reasoning process on LCB sample 259 under the MHC perturbation.
In §4.3.2, we identified a critical vulnerability in QwQ-32B’s internal reasoning. By analyzing failure cases that generated over 32,000 tokens under the MHC perturbation (occurring in 4 out of 479 tasks), we observed an unhealthy self-reflection loop and a progressive loss of confidence. We manually segmented the reasoning process into six stages (see Figure 4 and Appendix E.6). Initially, the model accurately computes the correct output, but the injected incorrect hint triggers it to repeatedly doubt its logic and re-execute the reasoning steps. This recursive verification gradually evolves into a brute-force enumeration, causing the model to generate hallucinated explanations. Ultimately, the reasoning process entirely collapses, with the model repeatedly generating over 12,000 “Hmm.”
To investigate the model’s confidence before collapsing, we track the cumulative frequency of confusing words and visualize the trend in Figure 2. The trend exhibits a striking quadratic growth pattern, indicating compounding confusion as reasoning progresses. The quadratic trajectory supports our stage-wise analysis: the model initially shows stable reasoning, but once the misleading hint is triggered, confusion accumulates, confidence undermines, culminating in complete failure.
@@ 6 Conclusion
We introduce a direction to evaluate the robustness of LLMs in code comprehension and execution under four structural and three textual perturbations. Our evaluation of seventeen LLMs under direct and CoT inference reveals critical vulnerabilities in model behavior, a heavy reliance on NL cues, and a limited capacity to follow actual program logic. While larger and newer models and CoT prompting significantly improve robustness against perturbations, they still cannot fully escape this surface-level dependence and leverage these cues to form reasoning shortcuts. Despite three LRMs demonstrating superior robustness under all perturbations—highlighting the strength of internal reasoning—we expose a new risk of overthinking and collapse when plausible but incorrect cues mislead their self-reflection mechanisms. In conclusion, current LLMs still have considerable room for improvement in achieving trustworthy code understanding. We believe that our benchmark serves as a practical framework for studying LLMs’ code inference behavior, providing significant insights into their robustness, reasoning tendencies, and potential failures.
Appendix A Illustrative Examples of Perturbation Strategies
Figure 3: Overview of CodeCrash pipeline. Code samples are first standardized into a clean vanilla version. Perturbations are visualized with yellow highlights. Each perturbed code is validated and executed before passing to LLMs. [Back To Execution Pipeline]
In section 3.3, we briefly describe all perturbation strategies employed in this paper. In this appendix section, we illustrate the detailed implementation for all perturbations with a LCB sample 36 as the example. Code 1 presents the base vanilla example where the formal answer is minimumCost(s = ’0011’) == 2.
@@ Appendix A Illustrative Examples of Perturbation Strategies
A.1 Renaming Entities
We use AST to identify all variable names appearing in:
@@ A.1 Renaming Entities
Code 2: REN-perturbed example from LCB (sample 36) for output prediction.
@@
A.2 Reformatting Conditional Expression
We generally categorize conditional expressions into three types: (1) Comparison-based conditions (e.g., if a < b:), (2) Boolean constant conditions (e.g., if True:), and (3) General conditions (e.g., if x:). For both comparison-based and boolean constant conditions, we design six transformation templates:
@@ A.2 Reformatting Conditional Expression
    bool(-~({cond})) == ({cond}),
@@
Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 3 presents an example of this perturbation.
Code 3: RTF-perturbed example from LCB (sample 36) for output prediction.
@@
A.3 Inserting Garbage Code
As introduced in §3.3.1, we inject three types of garbage code: (1) repeated-name global variable declarations, (2) in-executable conditional statements, and (3) dead-loop function definitions. For (1), we randomly select a few parameters or variables and declare global variables using the same names, assigning arbitrary integers at the beginning of the program. For (2), we design seven templates for false conditions:
@@ A.3 Inserting Garbage Code
Code 4: Garbage death loop function templates for GBC perturbation.
@@
Code 5: GBC-perturbed example from LCB (sample 36) for output prediction.
@@
A.4 Aggregating Structural Perturbations
As introduced in §3.3.1, we aggregate the application of REN, RTF, and GBC to a single code snippet, as shown in Code 6.
Code 6: ALL-perturbed example from LCB (sample 36) for output prediction.
Var_1 = 24
def f2(Var_1: str) -> int:
def f1():
for Var_4 in iter(lambda : True, False):
pass
def funct6():
for Var_3 in iter(int, 1):
Var_3 += 1
Var_6 = 0
for Var_3 in range(1, len(Var_1)):
Var_7 = Var_1 if 0 else Var_1
if None:
Var_5 = Var_1
if (lambda : Var_1[Var_3 - 1] != Var_1[Var_3])():
Var_6 += min(Var_3, len(Var_1) - Var_3)
while Var_1 != Var_1:
Var_2 = Var_1
return Var_6
A.5 Inserting Misleading Messages
As introduced in §3.3.2, we insert misleading comments for key AST node types, including function declarations, return statements, loops, conditionals, variable assignments, and common data structure operations. The data structure operations include common list, set, and string manipulations: append, extend, insert, remove, pop, sort, reverse, update, add, split, join, replace, lower, upper, capitalize, and swap case. To construct these comments, we design prompts for GPT-4o to generate misleading but plausible messages for each type. In this experiment, we manually selected 20 messages for each AST node type and 5 for each data structure operation. Take variable assignments as an example; Code 7 is the prompt for generating misleading comments for variable assignments, and Code 8 is a list of selected misleading comment candidates. Code 9 and 10 present an example of MDC and MPS perturbations.
@@ A.5 Inserting Misleading Messages
Code 8: Selected misleading comments for variable assignments used in MDC and MPS perturbations, where ‘{variable}’ is the variable name.
@@
Code 9: MDC-perturbed example from LCB (sample 36) for output prediction.
@@
A.6 Inserting Incorrect Hints
In order to generate plausible, logical but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression to incorrect ones and only allow it to modify the output value using the prompt in Code 11 for output prediction. An example perturbed instance is shown in Code 12, where the ground-truth return value 2 is replaced with the misleading hint # The return value is 3.
@@ A.6 Inserting Incorrect Hints
Appendix B Statistical Analysis

### v3

@@ 4.3 Defenses and Limitations
Table 8: Performance of models with and without explicit comment-ignoring prompts under MCC and MHC perturbations.
Model VAN MCC MHC
@@
DeepSeek-V3 (Direct) $67.8$ $39.7$ $41.8$ $47.7$ $47.8$
DeepSeek-V3 (CoT) $92.1$ $70.2$ $82.1$ $90.0$ $94.6$
GPT-4o (Direct) $64.5$ $45.7$ $45.9$ $47.3$ $46.8$
GPT-4o (CoT) $94.4$ $87.5$ $89.4$ $94.8$ $94.6$
Qwen2.5-72B-Instruct (Direct) $54.9$ $48.0$ $48.0$ $32.6$ $32.5$
Qwen2.5-72B-Instruct (CoT) $90.4$ $85.4$ $86.0$ $82.7$ $84.3$
Table 9: Performance of GPT-4o with and without refactoring under direct inference.
Perturbation w/o w/
VAN $64.5$ -
PSC-ALL $46.2$ $55.3$
MCC $\mathbf{45.7}$ $\mathbf{41.1}$
MPS $49.6$ $54.9$
Prompt engineering provides only partial relief.
@@
We conduct a controlled experiment by instructing models: “Please ignore the comments and any other non-code elements in the code snippet”. Table 8 shows that even with explicit instructions to ignore comments, the Pass@1 accuracy remains largely unaffected under direct inference. Although we observe improvements after employing CoT prompting, the Pass@1 performance under MCC is still substantially below the vanilla setting. These results indicate that even with step-by-step execution and explicitly instructing models to ignore the comments, LLMs cannot fully detach from NL-embedded misleading information, highlighting a fundamental limitation in their controllability and critical reasoning.
Refactoring agents improve syntax, but risk logic distortion.
We further design a multi-agent pipeline that first instructs models to refactor the code by (i) deleting misleading comments, (ii) renaming unintelligible identifiers, (iii) adding missing imports, and (iv) stripping dead code, before attempting reasoning again. Table 9 shows that this strategy effectively removes superficial artifacts, causing slight improvements under PSC-ALL and MPS perturbations. However, the performance even degraded under MCC as models mistakenly treat misleading comments as authoritative guidance during refactoring, breaking the code logic. Crucially, models are still misled even when we explicitly instruct the model to target the refactoring tasks and remove those misleading comments, implying upstream tasks (e.g., bug fixing and vulnerability detection) are equally at risk because an attacker can leverage benign-looking comments to poison the code and propagate failures. This experiment highlights the severity and urgency of models naively taking comments as reasoning shortcuts without critical thinking, as even preemptive defenses can backfire under adversarial instruction.
5 Related Work
Code generation and execution benchmarks.
Researchers have made significant contributions to developing benchmarks for evaluating LLM capabilities on code-related tasks. Early benchmarks such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021), APPS Hendrycks et al. (2021), and HumanEval+ Liu et al. (2023b) primarily focused on assessing code generation performance. More recent efforts have shifted toward practical software engineering and realistic coding scenarios. For instance, SWE-bench Jimenez et al. (2024), InfiBench Li et al. (2024), and LCB Jain et al. (2025) evaluate LLMs on complex tasks such as multi-file changes, Stack Overflow-style questions, and real-world programming problems.
Code execution and reasoning.
A growing body of work investigates LLMs in code execution. Early studies Austin et al. (2021); Nye et al. (2021) indicated that LLMs face challenges in accurately executing code but can benefit from intermediate reasoning steps. Subsequent research has utilized various strategies to enhance execution performance, including execution-informed pretraining Liu et al. (2023a), CoT prompting with trace supervision Ni et al. (2024), and iterative instruction prompting Lyu et al. (2024). Crux Gu et al. (2024) formalized the input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning. REVAL Chen et al. (2024b) and CACP Hooda et al. (2024) uncovered inconsistencies and concept-level misunderstandings of LLMs through intermediate reasoning analysis during code execution. CodeScore Dong et al. (2024), Lever Ni et al. (2023), and xCodeEval Khan et al. (2024) have leveraged code execution to improve model evaluation methodologies and enhance the performance of downstream tasks.
Robustness evaluation for code LLMs.
Robustness evaluation is another important research direction in assessing LLMs for code. Prior studies focused on prompt-level variations in code generation, examining how semantically equivalent NL descriptions impact model performance Mastropaolo et al. (2023); Chen et al. (2024a); Zhuo et al. (2023). Apart from NL perturbations, researchers have investigated model robustness under programming language perturbations. For example, CCTest Li et al. (2023) introduced structural mutations, CREAM Gao et al. (2023) employed counterfactual reasoning, and ReCode Wang et al. (2022) applied semantic-preserving transformations, advancing robustness in code completion. BigCodeBench Zhuo et al. (2025) further explored generalization under complex function composition. In addition, many works have explored structure-aware modeling by embedding program graphs or flow signals to enhance code robustness Son et al. (2022); Oh and Yoo (2024); Pei et al. (2022); Tipirneni et al. (2024). Furthermore, Chen et al. (2025a); Cuadron et al. (2025) and Kumar et al. (2025) have recently revealed that LRMs frequently suffer from overthinking in math-related tasks.
@@ 6 Conclusion
We introduce CodeCrash, a framework for stress-testing the robustness of LLMs in code reasoning under structural and NL-embedded perturbations. Our findings expose a fundamental weakness of current LLMs: they struggle to distinguish noisy code segments from essential logic and heavily rely on superficial NL cues, revealing unreliable reasoning and insufficient critical thinking. Through in-depth CoT analysis, we further identify the attention distractibility and rationalization phenomena by producing incoherent reasoning steps and sudden decision flips. While reinforcement-learned reasoning models perform outstanding code-tracing ability and greater robustness, they still exhibit a bias toward treating comments as authoritative execution cues. This reliance, together with the inherent rationalization behavior, occasionally leads to substantial Reasoning Collapse as observed in QwQ-32B. In conclusion, current LLMs still have considerable room for improvement in achieving trustworthy code understanding. We believe that our benchmark serves as a practical framework to provide significant insights into model robustness in code reasoning.
Acknowledgments and Disclosure of Funding
The paper was supported by two grants from the Research Grants Council of the Hong Kong Special Administrative Region, China: (1) No. CUHK 14209124 of the General Research Fund, and (2) No. SRFS2425-4S03 of the Senior Research Fellow Scheme.
Appendix A Dataset Selection
We evaluate two benchmarks: Crux Gu et al. [2024] and LCB Jain et al. [2025]. Crux contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input and output prediction tasks. In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on output prediction. While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computation.
Appendix B Illustrative Examples of Perturbation Strategies
In section 2.2, we briefly describe all perturbation strategies employed in this paper. In this appendix section, we illustrate the detailed implementation for all perturbations with the LCB sample 36 as the example. Code 1 presents the base vanilla example where the formal answer is minimumCost(s = ’0011’) == 2.
@@ Appendix B Illustrative Examples of Perturbation Strategies
B.1 Aggregating Structural Perturbations (PSC-ALL)
Table 10: Relative performance degradation under individual structural perturbations compared with the aggregated PSC-ALL perturbation.
Model Series Model Name VAN REN RTF GBC PSC-ALL Average
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.3$ $53.2$ $1.1\%$ $18.2\%$ $3.1\%$ $1.8\%$ $19.1\%$ $10.3\%$ $27.9\%$ $31.9\%$ $13.8\%$
GPT-4o $71.3$ $64.5$ $3.8\%$ $14.9\%$ $4.7\%$ $7.1\%$ $12.9\%$ $5.0\%$ $15.0\%$ $28.4\%$ $10.9\%$
Claude Series Claude-3.5-Haiku-20241022 $57.4$ $58.2$ $6.9\%$ $16.0\%$ $3.1\%$ $0.6\%$ $12.7\%$ $7.8\%$ $24.0\%$ $36.9\%$ $12.9\%$
Claude-3.5-Sonnet-20241022 $71.5$ $73.8$ $3\% 22.$$2.7\%$ $1.3\%$ $1.6\%$ $9.6\%$ $6.7\%$ $14.8\%$ $34.1\%$ $10.2\%$
Gemini Series Gemini-1.5-Flash $56.2$ $44.3$ $6.3\%$ $0.9\%$ $2.1\%$ $2.4\%$ $8.2\%$ $0.8\%$ $18.4\%$ $12.7\%$ $7.0\%$
Gemini-2.0-Flash $65.0$ $66.0$ $3.6\%$ $27.4\%$ $3.2\%$ $1.2\%$ $23.8\%$ $13.3\%$ $33.2\%$ $42.1\%$ $17.8\%$
Gemini-1.5-Pro-002 $67.4$ $56.2$ $4.3\%$ $6.1\%$ $3.7\%$ $8.4\%$ $8.3\%$ $6.8\%$ $19.9\%$ $23.2\%$ $9.8\%$
DeepSeek DeepSeek-V3 $67.9$ $67.8$ $3.9\%$ $24.1\%$ $2.0\%$ $5.5\%$ $6.8\%$ $8.8\%$ $12.9\%$ $35.6\%$ $10.9\%$
LLaMA Series LLaMA-3.1-8B-Instruct $36.0$ $34.7$ $0.2\%$ $12.4\%$ $0.9\%$ $4.8\%$ $8.7\%$ $13.7\%$ $23.8\%$ $19.7\%$ $9.9\%$
LLaMA-3.1-70B-Instruct $56.1$ $44.9$ $4.2\%$ $3.1\%$ $1.0\%$ $3.9\%$ $9.6\%$ $5.4\%$ $19.6\%$ $17.2\%$ $8.2\%$
LLaMA-3.1-405B-Instruct $63.5$ $50.7$ $8.5\%$ $7.7\%$ $1.0\%$ $8.8\%$ $9.1\%$ $5.6\%$ $19.7\%$ $16.9\%$ $9.6\%$
LLaMA-3.3-70B-Instruct $59.9$ $48.5$ $3.7\%$ $2.9\%$ $2.1\%$ $7.0\%$ $9.2\%$ $11.8\%$ $17.3\%$ $20.1\%$ $9.0\%$
Qwen Series Qwen2.5-7B-Instruct $43.3$ $41.4$ $7.0\%$ $17.5\%$ $3.0\%$ $7.7\%$ $29.9\%$ $9.1\%$ $37.9\%$ $30.9\%$ $18.3\%$
Qwen2.5-14B-Instruct $47.8$ $49.5$ $3.8\%$ $20.0\%$ $1.4\%$ $2.8\%$ $23.5\%$ $6.9\%$ $39.8\%$ $30.0\%$ $16.3\%$
Qwen2.5-32B-Instruct $60.0$ $59.6$ $2.8\%$ $18.1\%$ $4.3\%$ $0.8\%$ $9.6\%$ $2.3\%$ $19.8\%$ $34.9\%$ $11.0\%$
Qwen2.5-72B-Instruct $60.1$ $54.9$ $1.7\%$ $15.5\%$ $3.1\%$ $0.8\%$ $17.8\%$ $3.7\%$ $23.3\%$ $25.2\%$ $11.3\%$
Qwen2.5-Coder-32B-Instruct $67.0$ $56.6$ $3.7\%$ $9\% 1.4-1.4\%$ $3.6\%$ $16.1\%$ $4.9\%$ $22.3\%$ $27.9\%$ $11.6\%$
Average $4.0\%$ $14.3\%$ $2.4\%$ $3.7\%$ $13.8\%$ $7.2\%$ $22.9\%$ $27.5\%$
$\mathbf{-7.8}\%$ $\mathbf{-2.9}\%$ $\mathbf{-11.3}\%$ $\mathbf{-24.6}\%$ $\mathbf{-11.7}\%$
As illustrated in §2.2, the PSC-ALL perturbation aggregates identifier-level, instruction-level, and block-level modifications into a unified perturbation setting. Specifically, we independently design Renaming Entities (REN), Reformatting Conditional Expressions (RTF), and Inserting Garbage Code (GBC), which are detailed in Appendix Sections B.1.1, B.1.2, and B.1.3, respectively. Code 2 shows an example of an PSC-ALL-perturbed program.
Compounding Effect of PSC-ALL.
Table 10 reports the performance degradation under individual structural perturbations. Compared to Table 1, which highlights NL-embedded misleading perturbations, REN, RTF, and GBC alone exhibit mild effects. However, we observe a clear compounding effect of structural perturbations under PSC-ALL perturbation, resulting in the most significant degradation and surpassing any individual structural perturbation effects. Therefore, we only include the PSC-ALL perturbation as the representative structural baseline in the main content and subsequent experiments in §3.3, §3.5, and §3.4 for structural perturbations.
Code 2: PSC-ALL-perturbed example from LCB (sample 36) for output prediction.
Var_1 = 24
def f2(Var_1: str) -> int:
def f1():
for Var_4 in iter(lambda : True, False):
pass
def funct6():
for Var_3 in iter(int, 1):
Var_3 += 1
Var_6 = 0
for Var_3 in range(1, len(Var_1)):
Var_7 = Var_1 if 0 else Var_1
if None:
Var_5 = Var_1
if (lambda : Var_1[Var_3 - 1] != Var_1[Var_3])():
Var_6 += min(Var_3, len(Var_1) - Var_3)
while Var_1 != Var_1:
Var_2 = Var_1
return Var_6
B.1.1 Renaming Entities (REN)
We use the AST to identify all variable names appearing in:
@@ B.1.1 Renaming Entities (REN)
Code 3: REN-perturbed example from LCB (sample 36) for output prediction.
@@
B.1.2 Reformatting Conditional Expressions (RTF)
We categorize conditional expressions into three general types: (1) comparison-based conditions (e.g., if a < b:), (2) boolean constant conditions (e.g., if True:), and (3) general conditions (e.g., if x:). For both comparison-based and Boolean constant conditions, we design seven transformation templates:
@@ B.1.2 Reformatting Conditional Expressions (RTF)
    bool(-˜({cond})) == ({cond}),
@@
Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 4 presents an example of this perturbation.
Code 4: RTF-perturbed example from LCB (sample 36) for output prediction.
@@
B.1.3 Inserting Garbage Code (GBC)
We inject three types of garbage code: (1) repeated-name global variable declarations, (2) unexecutable conditional statements, and (3) dead-loop function definitions. For (1), we randomly select a few parameters or variables and declare global variables using the same names, assigning arbitrary integers at the beginning of the program. For (2), we design seven templates for false conditions:
@@ B.1.3 Inserting Garbage Code (GBC)
Code 5: Garbage death loop function templates for GBC perturbation.
@@
Code 6: GBC-perturbed example from LCB (sample 36) for output prediction.
@@
B.2 Inserting Contextual-Level Misleading Messages (MCC and MPS)
As introduced in §2.2, we insert misleading comments for eight key AST node types, including (1) function declarations, (2) return statements, (3-4) while and for loops, (5) conditionals, (6) variable assignments, and (7) common data structure operations. The (8) data structure operations include common list, set, and string manipulations: append, extend, insert, remove, pop, sort, reverse, update, add, split, join, replace, lower, upper, capitalize, and swap case. To construct these comments, we design prompts for GPT-4o to generate misleading but plausible messages for each type. In this experiment, we manually selected 20 messages for each AST node type and 5 for each data structure operation. Take variable assignments as an example; Code 7 is the prompt for generating misleading comments for variable assignments, and Code 8 is a list of selected misleading comment candidates. Code 9 and 10 present an example of MCC and MPS perturbations.
@@ B.2 Inserting Contextual-Level Misleading Messages (MCC and MPS)
Code 8: Selected misleading comments for variable assignments used in MCC and MPS perturbations, where ‘{variable}’ is the variable name.
@@
Code 9: MCC-perturbed example from LCB (sample 36) for output prediction.
@@
B.3 Inserting Reasoning-Level Misleading Hint (MHC)
In order to generate plausible, logical, but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression into incorrect ones and only allow it to modify the output value using the prompt in Code 11 for output prediction. An example perturbed instance is shown in Code 12, where the ground-truth return value 2 is replaced with the misleading hint # The return value is 3.
@@ B.3 Inserting Reasoning-Level Misleading Hint (MHC)
Appendix C Statistical Analysis